# AI4EOSC: a Federated Cloud Platform for Artificial Intelligence in Scientific Research

### 17 Mar 2026 v2

@@ 5.2.3 AI4Life
AI4Life [^9] https://ai4life.eurobioimaging.eu/, a Horizon Europe funded project to empower life science researchers to harness the full potential of AI/ML methods for bioimage analysis, began collaborating with AI4EOSC in order to leverage the AI4EOSC software stack and platform. The AI4EOSC and AI4Life teams worked on both platforms integration, as a means to implement and pilot a model for interoperability across research infrastructures, aligned with FAIR principles. This integration enables efficient sharing and use of AI/ML models and computing resources across domains.
@@ 6 Conclusion and Future Work
With a fragmented AI ecosystem increasingly dominated by large corporations, a scalable open-source solution putting the focus on reproducibility and interoperability is needed. To address this challenge, in the context of science and research, we have here presented the AI4EOSC platform. We provide a comprehensive set of tools that target the whole AI/ML lifecycle, backed with easy to use interfaces and dashboards, a federated workload management system to train models, options to deploy inference models along the whole Cloud Continuum, the ability to create composable inference workflows paired with our AI/ML as a Service serverless framework, built-in AI model provenance to foster trustworthiness, and an extensive CI/CD pipeline that ensures AI code quality and enables MLOps practices.
The AI4EOSC architecture provides a modular and extensible reference that allows to build upon it, enabling the integration of external services and catalogs. It provides an overall customization system that has been adopted by relevant external research communities, and that is enabling and advancing FAIR development of AI/ML in science and within the EOSC. In these lines, we are working with specific user communities in order to integrate their community tools (e.g., data repositories, virtual research environments) with the platform.
Concerning future work, we plan to keep evolving the platform along several axes. Specifically, we plan to perform extensive work in interoperability, and namely we will extend the semantic support of ML assets, like datasets or models, to be able to seamlessly import, export and connect them in the platform, fostering semantic interoperability following open standards such as Croissanl-ML [2]. In addition, we plan to extend the usage of the Experiment Tracking component to serve as a model registry, enriching the models with semantic information. We will also further strengthen component integration, for example, by allowing the use of any existing AI4EOSC AI module as an annotation assistant inside the CVAT tool, using the AIaaS serverless endpoints.
Besides semantic interoperability, we plan to work on the technical interoperability. More specifically, we plan to implement the OAI-PMH protocol at the marketplace level, making it possible to automatically harvest the module metadata. Additional integrations and interoperability with external marketplaces and model repositories (both science related and generalistic) are in our roadmap.
Lastly, we are also working towards making the platform aware of the energy consumption, both at the control plane and the user workloads. This will allow us to calculate the carbon and water footprint, integrating fine-grained and accurate impact measurements from Wattnet [18], and exposing them to the users, via the module metadata as well as upon workload submission.

### 05 May 2026 v3

@@ 6.2.3 AI4Life
AI4Life [^9] https://ai4life.eurobioimaging.eu/, a Horizon Europe funded project to empower life science researchers to harness the full potential of AI/ML methods for bio-image analysis, began collaborating with AI4EOSC in order to leverage the AI4EOSC software stack and platform. The AI4EOSC and AI4Life teams worked on both platforms integration, as a means to implement and pilot a model for interoperability across research infrastructures, aligned with FAIR principles. This integration enables efficient sharing and use of AI/ML models and computing resources across domains.
@@ 7 Conclusion and Future Work
The growing adoption of AI and ML in scientific research demands platforms that go beyond industrial MLOps frameworks to address the specific needs of open science: reproducibility, FAIR compliance, provenance tracking, and seamless integration with federated research infrastructures to enable easy access to the technologies by researchers. This paper has presented AI4EOSC, a federated, open-source platform that operationalizes these principles across the full AI/ML lifecycle within the European Open Science Cloud ecosystem.
The main architectural contribution of AI4EOSC is not the provision of any single or standalone tool, but rather the integration of development, training, deployment, and provenance into a unified, FAIR-by-design system that enforces open science compliance at the infrastructure level, alleviating the burden from individual researchers. AI4EOSC main innovations include a lightweight inter-cluster federation based on Consul and Nomad that avoids the infrastructure fragmentation and silos of Kubernetes-centric stacks, an automated W3C PROV-compliant provenance pipeline that captures full model lineage with limited user intervention, and extended federated learning support with token-based authentication, novel aggregation strategies, and server-side differential privacy. AI4EOSC value proposition has been validated through three real-world community adoptions: iMagine, KMD4EOSC, and AI4Life. This demonstrates consistent functionality and deployment across heterogeneous resource providers and diverse scientific domains.
Nevertheless several limitations need to be acknowledged. Firstly, the current provenance model, while comprehensive for single-model workflows, requires further work to handle multi-model inference pipelines and federated learning scenarios where data remains siloed across institutions. Secondly, although the platform currently supports semantic interoperability through MLDCAT-AP serialization, full machine-actionability of AI/ML assets across heterogeneous repositories and data spaces remains an open and evolving challenge: deeper integration with emerging standards and external catalogs is needed to allow assets to be not only findable and accessible, but autonomously actionable by external systems. Lastly, the operational overhead of maintaining a geographically distributed federation (including failure handling, performance variability across heterogeneous providers, and long-term governance) represents a practical challenge that will grow as the federation expands. Although automation is in place in order to alleviate this fact, this is a common issue of any geographically and complex distributed infrastructure, and AI4EOSC alleviates this fact by leveraging a single and unified control plane.
Future work in the near term will be focused on deepening semantic interoperability and machine actionability by extending the current support for MLDCAT-AP, and focusing on emerging standards such as Croissant-ML [3]. Moreover, we will implement the OAI-PMH protocol to enable automated metadata harvesting across external marketplaces and repositories. We will also extend the MLflow experiment tracking component to serve as a full model registry, enriching stored models with semantic provenance information, as we are already doing with the modules published through the dashboard. In the medium term, we plan to address the provenance gaps identified above, in particular for federated and multi-model scenarios, and to strengthen the platform’s energy and sustainability awareness by integrating fine-grained carbon and water footprint measurements via Wattnet [15], implementing platform level optimizations, and informing users about the impact of their workloads. In the longer term, we aim to grow the current federation by lowering the barrier for new resource providers to join, and to conduct systematic user studies to measure the actual impact of the platform on researcher productivity and reproducibility outcomes.